OpenStudy (anonymous):

determine the area of ABC with vertices A(-4,1,1) B(-2,0,3) C(-7,5,8)

OpenStudy (dan815):

let B^=B-A C^=C-a these are 2 vectors The magnitude of the crossproduct gives us their area

OpenStudy (anonymous):

im doing tht problem on flvs rite now thts weired

OpenStudy (anonymous):

could u show me